# Electricity Price Forecasting: Bridging Linear Models, Neural Networks and Online Learning

### v1

@@ Abstract
Precise day-ahead forecasts for electricity prices are crucial to ensure efficient portfolio management, support strategic decision-making for power plant operations, enable efficient battery storage optimization, and facilitate demand response planning. However, developing an accurate prediction model is highly challenging in an uncertain and volatile market environment. For instance, although linear models generally exhibit competitive performance in predicting electricity prices with minimal computational requirements, they fail to capture relevant nonlinear relationships. Nonlinear models, on the other hand, can improve forecasting accuracy with a surge in computational costs. We propose a novel multivariate neural network approach that combines linear and nonlinear feed-forward neural structures. Unlike previous hybrid models, our approach integrates online learning and forecast combination for efficient training and accuracy improvement. It also incorporates all relevant characteristics, particularly the fundamental relationships arising from wind and solar generation, electricity demand patterns, related energy fuel and carbon markets, in addition to autoregressive dynamics and calendar effects. Compared to the current state-of-the-art benchmark models, the proposed forecasting method significantly reduces computational cost while delivering superior forecasting accuracy (12-13% RMSE and 15-18% MAE reductions). Our results are derived from a six-year forecasting study conducted on major European electricity markets.
@@ 1 Introduction
The day-ahead forecast models can be classified into two main categories: Linear models and non-linear ones. Linear models with an autoregressive structure, where the lags of the price are used as predictors, were popular ([contreras2003arima]; [zhou2006electricity]). Later, the inclusion of exogenous variables such as solar and load along the lags has improved accuracy ([conejo2005forecasting]). More recently, shrinkage estimation methods such as least absolute shrinkage and selection operator (LASSO) ([tibshirani1996regression]) and Elastic Net ([zou2005regularization]) have demonstrated strong predictive accuracy thanks to their ability to select relevant features ([lear]; [ziel2016forecasting], [ziel2018day]). Linear models are generally effective in predicting electricity prices and are highly valued for their interpretability and fast computation. However, due to their basic structure, they may still miss important non-linear relationships, such as the interaction between variables which can lead to less accurate forecasts.
@@
Those interactions, on the other hand, could be captured by non-linear models, which generally consist of many classes, particularly from machine learning models such as decision tree learners, like random forests, and gradient boosting machines; and neural networks. However, we will focus here on the deep learning methods, given their rapidly growing popularity in electricity price forecasting ([a]). The multilayer perceptron (MLP) is one of the simple architectures that has already been employed to generate forecasts a few years back ([szkuta2002electricity]; [catalao2007short]). Later, a more sophisticated architecture, like a denoising autoencoder, originally popular in natural language processing, is employed for price forecasting ([wang2016short]). Additionally, Recurrent neural network (RNN) is also widely employed due to its structure to capture temporal dependencies in time series data ([dnn]; [ugurlu2018electricity]; [chinnathambi2018deep]). Note that the deep learning methods usually outperform the linear models in terms of accuracy but fails in terms of computational cost ([a]). To address this trade-off, we build a hybrid model that combines the strengths of both approaches while following the standard framework checklist proposed by [a].
@@
We propose a simple neural network architecture that combines input and output through two components. The first component is a generalized linear model that connects the input to the output directly, while the second is an MLP that connects the input to the output through one hidden layer with a nonlinear activation function. This architecture was first introduced by [google] to improve application recommendations. Thanks to this structure, the model can both learn the explicit linear relationship and the hidden nonlinear interactions. In addition to this unique architecture, the model incorporates all key drivers of electricity price formation, including renewable generation, electricity demand, fuel and carbon prices, autoregressive dynamics, and calendar effects. This ensures that both fundamental and seasonal relationships are fully captured. It is worth emphasizing that the literature is rich in hybrid models that combine different algorithms to improve accuracy in electricity price forecasting ([kim2015short]; [ben2018forecasting]; [bento2018bat]; [bisoi2020short]; [gomez2023electricity]; [cu2025time] ). However, none of the existing studies have employed the proposed hybrid model.
@@
To further improve accuracy, an ensemble method is implemented to average the individual forecast models. Combining the predictions of different models has been shown to enhance precision as compared to individual forecasts in many studies ([raviv2015forecasting], [nowotarski2016combine], [maciejowska2020pca]). The arithmetic mean is the simplest averaging technique, such that equal weight is assigned to each individual forecast ([stock2004combination]). In this work, we select fully adaptive Bernstein Online Aggregation (BOA) proposed by [wintenberger2017optimal] as an ensemble method. This approach assigns dynamic weights to individual models based on their recent performance. While fully adaptive BOA was previously used to predict the electricity load ([de2023adaptive]; [hirsch2024online]; [himych2024adaptive]), to the best of our knowledge, it has not yet been used for electricity price forecasting. .
@@
In addition to improving accuracy, keeping the computational time at an efficient level is a key objective of this work. To this end, we implement a novel partial online learning approach designed to minimize the computational burden. This method employs information from the previous iteration to reduce the training time of the model. While online learning algorithms were formerly employed in the electricity price literature, for instance [hirsch2025online] used it to generate a probabilistic price forecast, there is no paper in the literature that has applied the non-parametric online learning that we introduced to forecast electricity prices.
@@
  2. 2.
    We are the first to apply a fully adaptive BOA framework for model ensembling and forecast combination.
@@
    We introduce a new non-parametric partial online learning approach suitable for both linear and neural network models, enabling fast and efficient learning.
@@ 2 Data
The datasets used in this project are mainly extracted from the website https://www.entsoe.eu of the European Network of Transmission System Operators for Electricity (ENTSO-E), which is a platform that provides complete information on the European electricity market. We consider data from 2018-01-10 to 2025-01-14 for the two largest European electricity markets: German-Luxembourg [^1] Note that in October 2018 the Austrian-German-Luxembourg price zone got split. and Spain. Thus, each of the datasets contains 2298 days of hourly observations (55153 hours) of the main variables for the electricity market.
For each of the considered datasets, the objective variable is the hourly day-ahead electricity price, which we denote $P_{d,h}$ for the price at day $d$ and hour $h$ where $h=0,\ldots,23$. We also consider typical external data, particularly day-ahead forecasts from renewable energy sources, solar ($\text{Solar}_{d,h}$), wind onshore ($\text{WindOn}_{d,h}$) and wind offshore ($\text{WindOff}_{d,h}$), as well as day-ahead load forecasts ($\text{Load}_{d,h}$). Note that wind offshore is only available on the German-Luxembourg dataset but not in the Spanish one, as Spain currently has no large operational offshore wind generation facilities ([cerda2025present]). In addition to these fundamental variables, both datasets incorporate four electricity market-related commodity price time series. These are closing prices of short-term commodity futures, of TTF natural gas price ($\text{NGas}_{d}$), Brent oil price ($\text{Oil}_{d}$), Rotterdam Coal price ($\text{Coal}_{d}$) and carbon emission prices by the European Emission Allowances ($\text{EUA}_{d}$). The commodity prices are collected from refinitiv datastream ([refinitiv_datastream]).
@@ 3 Model Design
Now we proposes and compares various models designed explicitly for day-ahead electricity price forecasting models.We examine two main configurations, linear and combined structure. Afterwards we explain the proposed ensembling method in detail.
@@ 3.1 Linear Networks
This architecture consists of an input layer and an output layer with 24 neurons, where each neuron represents the price of electricity at a specific hour. This setup is employed as the basis for two variants, referred to as the Reduced Linear and Full Linear models. The Reduced Linear predicts the hourly electricity prices using regressors of the specific predicted hour. For instance, to predict hour zero (midnight to one o’clock in the morning), the model uses only the lagged prices and fundamentals corresponding to hour 0, along with time-invariant variables such as weekday dummies and commodity prices. The Reduced Linear model is expressed as follows:
$P_{d,h}$ $=\beta_{h,0}+\beta_{h,1}P_{d-1,h}+\underbrace{\beta_{h,2_{2}}P_{d-2,h}+\beta_{h,2_{3}}P_{d-7,h}}_{\text{$=\bm{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,h}$}}+\beta_{h,3}P_{d-1,23}+$
$\quad\underbrace{\beta_{h,4_{1}}\text{Solar}_{d,h}+\beta_{h,4_{2}}\text{WindOn}_{d,h}+\beta_{h,4_{3}}\text{WindOff}_{d,h}+\beta_{h,4_{4}}\text{Load}_{d,h}}_{\text{$=\bm{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,h}$}}+$
$\quad\underbrace{\beta_{h,5_{1}}\text{Oil}_{5-2}+\beta_{h,5_{2}}\text{Coal}_{d-2}+\beta_{h,5_{3}}\text{EUA}_{d-2}+\beta_{h,5_{4}}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\quad\underbrace{\beta_{h,6_{1}}\text{Mon}_{d}+\beta_{h,6_{2}}\text{Sat}_{d}+\beta_{h,6_{3}}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\epsilon_{d,h}$ (1)
@@
and the error term $\epsilon_{d,h}$. The coefficients $\beta_{d,j}$’s represent the model parameters to be estimated or trained. Commodity prices are modeled using information available on day d2d-2, since at the time of bidding (before noon on day d1d-1) the stock closing prices of commodities are not yet available.
In addition to the Reduced Linear model, we consider a Full Linear model, which uses regressors from all hours rather than restricting them to the target predicted hour $h$. For example, when predicting the price for hour $h=0$, the model incorporates lagged prices and fundamentals from hours 0 through 23 as predictors, together with regressors that do not vary within the day. The following equation provides the Full Linear model specification:
$P_{d,h}$ $=\beta_{h,0}+\sum_{l=0}^{23}\beta_{h,l,1}P_{d-1,l}+\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,2_{2}}P_{d-2,l}+\beta_{h,l,2_{3}}P_{d-7,l}\right)}_{\text{$=\bm{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,l}$}}+$
$\qquad\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,4_{1}}\text{Solar}_{d,l}+\beta_{h,l,4_{2}}\text{WindOn}_{d,l}+\beta_{h,l,4_{3}}\text{WindOff}_{d,l}+\beta_{h,l,4_{4}}\text{Load}_{d,l}\right)}_{\text{$=\bm{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,l}$}}+$
$\qquad\underbrace{\beta_{h,5_{1}}\text{Oil}_{d-2}+\beta_{h,5_{2}}\text{Coal}_{d-2}+\beta_{h,5_{3}}\text{EUA}_{d-2}+\beta_{h,5_{4}}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\qquad\underbrace{\beta_{h,6_{1}}\text{Mon}_{d}+\beta_{h,6_{2}}\text{Sat}_{d}+\beta_{h,6_{3}}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\nu_{d,h},$ (2)
@@
3.2 Non-Linear Networks and Combined structures
The second main architecture in this project is a more elaborate network that, in addition to input to the output layer, employs a hidden layer equipped with the Leaky Relu activation function to capture potential nonlinearities. Two specifications of this architecture are considered: the MLP with Reduced Linear and the MLP with Full Linear. The MLP with Reduced Linear is the sum of the linear predictor of equation (1) and a nonlinear part $\mathrm{NN}(\theta)$:
@@ 3.2 Non-Linear Networks and Combined structures
where $\mathbf{x}\in\mathbb{R}^{m}$ is the input vector with $m$ regressors, $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\mathbf{b}^{(1)},\mathbf{b}^{(2)}$ are the trainable weight matrices and bias vectors, respectively. $\mathbf{\theta}$ denotes the collection of all trainable parameters, and $\alpha=0.01$, is the leakage factor in the Leaky ReLU activation function.
Figure 3 illustrates the architecture of MLP with Reduced Linear Model. The green neurons represent the input layer. The hidden layer is defined by the blue $n$ neurons, where the number of neurons $n$ is a hyperparameter to be tuned. The output layer consists of 24 neurons, where each neuron provides the forecast for a specific hour of the day, is highlighted in orange. The blue edges linking the input and hidden layers correspond to the weight matrix $\mathbf{W}^{(1)}$, while the remaining blue connections are saved in $\mathbf{W}^{(2)}$. For simplicity, the bias vectors $\mathbf{b}^{(1)}$ and $\mathbf{b}^{(2)}$ are not shown in Figure 3. Note that the rectangular neurons represent vectors rather than scalars. Thus, any connection from a rectangular node to another neuron does not represent a single weight, but a vector of weights. For instance, the blue edges connecting $\mathbf{Com}_{d-2}$ to the hidden neuron $\mathbf{H}_{1}$ illustrate four learnable weights, since $\mathbf{Com}_{d-2}$ is a 4-dimensional vector.
Figure 3: MLP with Reduced Linear Model architecture to forecast the hourly electricity price
The red edges directly linking the input to the output, referred to as skip connection, represents the model’s linear predictor. The red solid lines connect to the variables that vary for each hour, whereas the dashed red lines correspond to variables that are constant throughout the day, such as fundamentals. However, an exception arises for the covariate $P_{d-1,23}$, which is constant across hours and also serves as the lagged price for predicting $P_{d,23}$. Accordingly, the skip connection for hours $P_{d,0}$ through $P_{d,22}$ includes 15 regressors, while the skip connection for $P_{d,23}$ contains only 14 covariates. It is worth noting that the linear relationships are not only learned through skip connection since the MLP component of the model can learn linear dependencies as well, which results in identifiability issues. However, one can include the orthogonalization cell proposed by [rugamer] to restore the interpretability of the linear components.
MLP with Full Linear model shares the same hidden layer design as the MLP with Reduced Linear model, which is defined as $\mathrm{NN}(\mathbf{\theta})$ in equation (3). Yet differs in how the input-to-output connection is structured, as this linear component corresponds to that of the Full Linear model. The model equation is expressed in the following form:
@@
Hence, the MLP with Full Linear model employs 175 regressors to forecast each hour.
In addition to the four models introduced above, we develop three further variants. The first is the plain MLP model, which consists of an input layer, a hidden layer, and an output layer, without any skip connections such that the coefficients of the linear component in equation (3) are set to zero ($\beta_{h,0}=\beta_{h,1}=\beta_{h,3}=\mathbf{\beta}_{h,4}^{\prime}=\mathbf{\beta}_{h,5}^{\prime}=\mathbf{\beta}_{h,6}^{\prime}=0$). The second and third are the Reduced Linear with OLS and the MLP with Reduced Linear and OLS models. These are structurally similar to the Reduced Linear and the MLP with Reduced Linear models, respectively, but differ in their weight initialization: the OLS-based variants use the coefficients of a simple linear model estimated via OLS as initial weights, whereas the non-OLS counterparts initialize weights randomly.
It is important to highlight that each of the above models is implemented identically to predict electricity prices in the German-Luxembourg and Spain bidding zones. The only exception is that $\beta_{h,4_{3}}$ is set to zero in the Spanish case due to the absence of offshore wind generation data. Furthermore, all the regressors are standardized to enhance training stability.

### v2

@@ Abstract
Precise day-ahead forecasts for electricity prices are crucial to ensure efficient portfolio management, support strategic decision-making for power plant operations, enable efficient battery storage optimization, and facilitate demand response planning. However, developing an accurate prediction model is highly challenging in an uncertain and volatile market environment. For instance, although linear models generally exhibit competitive performance in predicting electricity prices with minimal computational requirements, they fail to capture relevant nonlinear relationships. Nonlinear models, on the other hand, can improve forecasting accuracy with a surge in computational costs. We introduce a novel partial online learning approach, the key contribution of this work, which substantially reduces computational time. In addition, we propose a multivariate hybrid neural architecture that combines linear and nonlinear feed-forward neural structures. Unlike previous hybrid models, our approach integrates forecast combination using Bernstein Online Aggregation (BOA) to further improve forecasting accuracy. Compared to the current state-of-the-art benchmark models, the proposed forecasting method significantly reduces computational cost while delivering superior forecasting accuracy (11-12% RMSE and 14-17% MAE reductions). Our results are derived from a six-year forecasting study conducted on major European electricity markets.
@@ 1 Introduction
The day-ahead forecast models can be classified into two main categories: Linear models and nonlinear ones. Linear models with an autoregressive structure, where the lags of the price are used as predictors, were popular ([contreras2003arima]; [zhou2006electricity]). Later, the inclusion of exogenous variables such as solar and load along the lags has improved accuracy ([conejo2005forecasting]). More recently, shrinkage estimation methods such as least absolute shrinkage and selection operator (LASSO) ([tibshirani1996regression]) and Elastic Net ([zou2005regularization]) have demonstrated strong predictive accuracy thanks to their ability to select relevant features ([lear]; [ziel2016forecasting], [ziel2018day]). Linear models are generally effective in predicting electricity prices and are highly valued for their interpretability and fast computation. However, due to their basic structure, they may still miss important nonlinear relationships, such as the interaction between variables which can lead to less accurate forecasts.
@@
Those interactions, on the other hand, could be captured by nonlinear models, which generally consist of many classes, particularly from machine learning models such as decision tree learners, like random forests, and gradient boosting machines; and neural networks. However, we will focus here on the deep learning methods, given their rapidly growing popularity in electricity price forecasting ([a]). The multilayer perceptron (MLP) is one of the simple architectures that has already been employed to generate forecasts a few years back ([szkuta2002electricity]; [catalao2007short]). Later, a more sophisticated architecture, like a denoising autoencoder, originally popular in natural language processing, is employed for price forecasting ([wang2016short]). Additionally, Recurrent neural network (RNN) is also widely employed due to its structure to capture temporal dependencies in time series data ([dnn]; [ugurlu2018electricity]; [chinnathambi2018deep]). Note that the deep learning methods usually outperform the linear models in terms of accuracy but fails in terms of computational cost ([a]). To address this trade-off, we build a hybrid model that combines the strengths of both approaches while following the standard framework checklist proposed by [a].
@@
We propose a simple neural network architecture that combines input and output through two components. The first component is a generalized linear model that connects the input to the output directly, while the second is an MLP that connects the input to the output through one hidden layer with a nonlinear activation function. This architecture was first introduced by [google] to improve application recommendations. Thanks to this structure, the model explicitly captures the linear structure commonly present in electricity price forecasting. Consequently, the nonlinear component only needs to model the remaining nonlinear patterns, which in turn may restrict the effective parameter search region for the nonlinear network and potentially reduce estimation error. In addition to this unique architecture, the model incorporates all key drivers of electricity price formation, including renewable generation, electricity demand, fuel and carbon prices, autoregressive dynamics, and calendar effects. This ensures that both fundamental and seasonal relationships are fully captured. It is worth emphasizing that the literature is rich in hybrid models that combine different algorithms to improve accuracy in electricity price forecasting ([kim2015short]; [ben2018forecasting]; [bento2018bat]; [bisoi2020short]; [gomez2023electricity]; [cu2025time] ). However, none of the existing studies have employed the proposed hybrid model.
@@
To further improve accuracy, an ensemble method is implemented to average the individual forecast models. Combining the predictions of different models has been shown to enhance precision as compared to individual forecasts in many studies ([raviv2015forecasting], [nowotarski2016combine], [maciejowska2020pca]). The arithmetic mean is the simplest averaging technique, such that equal weight is assigned to each individual forecast ([stock2004combination]). In this work, we select fully adaptive Bernstein Online Aggregation (BOA) proposed by [wintenberger2017optimal] as an ensemble method. This approach assigns dynamic weights to individual models based on their recent performance. While fully adaptive BOA was previously used to predict the electricity load ([de2023adaptive]; [hirsch2024online]; [himych2024adaptive]), to the best of our knowledge, it has not yet been used for electricity price forecasting.
@@
In addition to improving accuracy, keeping the computational time at an efficient level is a key objective of this work. To this end, we implement a novel partial online learning approach designed to minimize the computational burden. This method integrates warm starting with window-specific hyperparameter configurations to reduce the training time of the model. Online learning algorithms were formerly employed in the electricity price literature, for instance [hirsch2025online] used it to generate a probabilistic price forecast. However, to the best of our knowledge, no study in the energy market literature or beyond has applied the proposed partial online learning approach.
@@
    We introduce a new partial online learning approach suitable for both linear and neural network models, enabling fast and efficient learning.
  2. 2.
@@
    We are the first to apply a fully adaptive BOA framework for model ensembling and forecast combination in electricity price forecasting.
@@ 2 Data
The datasets used in this project are mainly extracted from the website https://www.entsoe.eu of the European Network of Transmission System Operators for Electricity (ENTSO-E), which is a platform that provides complete information on the European electricity market. We consider data from 2018-10-01 to 2025-01-14 for the two largest European electricity markets: German-Luxembourg [^1] Note that in October 2018 the Austrian-German-Luxembourg price zone got split. and Spain. Thus, each of the datasets contains 2298 days of hourly observations (55153 hours) of the main variables for the electricity market.
For each of the considered datasets, the objective variable is the hourly day-ahead electricity price, which we denote $P_{d,h}$ for the price at day $d$ and hour $h$ where $h=0,\ldots,23$. We also consider typical external data, particularly day-ahead forecasts from renewable energy sources, solar ($\text{Solar}_{d,h}$), wind onshore ($\text{WindOn}_{d,h}$) and wind offshore ($\text{WindOff}_{d,h}$), as well as day-ahead load forecasts ($\text{Load}_{d,h}$). Note that wind offshore is only available on the German-Luxembourg dataset but not in the Spanish one, as Spain has no large operational offshore wind generation facilities ([cerda2025present]). In addition to these fundamental variables, both datasets incorporate four electricity market-related commodity price time series. These are closing prices of short-term commodity futures, of TTF natural gas price ($\text{NGas}_{d}$), Brent oil price ($\text{Oil}_{d}$), Rotterdam Coal price ($\text{Coal}_{d}$) and carbon emission prices by the European Emission Allowances ($\text{EUA}_{d}$). The commodity prices are collected from refinitiv datastream ([refinitiv_datastream]).
@@ 3 Model Design
Now we propose and compare various models designed explicitly for day-ahead electricity price forecasting models. We examine two main configurations, linear and combined structure. Afterwards we explain the proposed ensembling method in detail.
@@ 3.1 Linear Networks
This architecture consists of an input layer and an output layer with 24 neurons, where each neuron represents the price of electricity at a specific hour. This setup is employed as the basis for two variants, referred to as the Reduced Linear (RLin) and Full Liner (FLin) models. The RLin model predicts the hourly electricity prices using regressors of the specific predicted hour. For instance, to predict hour zero (midnight to one o’clock in the morning), the model uses only the lagged prices and fundamentals corresponding to hour 0, along with time-invariant variables such as weekday dummies and commodity prices. The RLin model is expressed as follows:
$P_{d,h}$ $=\beta_{h,0}+\beta_{h,1}P_{d-1,h}+\underbrace{\beta_{h,2,1}P_{d-2,h}+\beta_{h,2,2}P_{d-7,h}}_{\text{$=\bm{\beta}_{h,2}^{\prime}\mathbf{P}_{\cdot,h}$}}+\beta_{h,3}P_{d-1,23}+$
$\quad\underbrace{\beta_{h,4,1}\text{Solar}_{d,h}+\beta_{h,4,2}\text{WindOn}_{d,h}+\beta_{h,4,3}\text{WindOff}_{d,h}+\beta_{h,4,4}\text{Load}_{d,h}}_{\text{$=\bm{\beta}_{h,4}^{\prime}\mathbf{Fun}_{d,h}$}}+$
$\quad\underbrace{\beta_{h,5,1}\text{Oil}_{d-2}+\beta_{h,5,2}\text{Coal}_{d-2}+\beta_{h,5,3}\text{EUA}_{d-2}+\beta_{h,5,4}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\quad\underbrace{\beta_{h,6,1}\text{Mon}_{d}+\beta_{h,6,2}\text{Sat}_{d}+\beta_{h,6,3}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\epsilon_{d,h}$ (1)
@@
  • $\epsilon_{d,h}$ is the error term.
The coefficients $\beta_{d,j}$’s represent the model parameters to be estimated or trained. Commodity prices are modeled using information available on day d2d-2, since at the time of bidding (before noon on day d1d-1) the stock closing prices of commodities are not yet available. Lagged prices, fundamental variables, commodity prices, and calendar variables are selected as regressors as they are the most well-known predictors in the electricity price forecasting literature. While $P_{d-1,23}$ is included based on empirical evidence in [ziel2016forecasting] that the last hour of the previous day strongly influences early-day electricity prices.
In addition to the RLin model, we consider a FLin model, which uses regressors from all hours rather than restricting them to the target predicted hour $h$. For example, when predicting the price for hour $h=0$, the model incorporates lagged prices and fundamentals from hours 0 through 23 as predictors, together with regressors that do not vary within the day. The following equation provides the FLin model specification:
$P_{d,h}$ $=\beta_{h,0}+\sum_{l=0}^{23}\beta_{h,l,1}P_{d-1,l}+\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,2,1}P_{d-2,l}+\beta_{h,l,2,2}P_{d-7,l}\right)}_{\text{$=\bm{\beta}_{h,l,2}^{\prime}\mathbf{P}_{\cdot,l}$}}+$
$\qquad\sum_{l=0}^{23}\underbrace{\left(\beta_{h,l,4,1}\text{Solar}_{d,l}+\beta_{h,l,4,2}\text{WindOn}_{d,l}+\beta_{h,l,4,3}\text{WindOff}_{d,l}+\beta_{h,l,4,4}\text{Load}_{d,l}\right)}_{\text{$=\bm{\beta}_{h,l,4}^{\prime}\mathbf{Fun}_{d,l}$}}+$
$\qquad\underbrace{\beta_{h,5,1}\text{Oil}_{d-2}+\beta_{h,5,2}\text{Coal}_{d-2}+\beta_{h,5,3}\text{EUA}_{d-2}+\beta_{h,5,4}\text{NGas}_{d-2}}_{\text{$=\bm{\beta}_{h,5}^{\prime}\mathbf{Com}_{d-2}$}}+$
$\qquad\underbrace{\beta_{h,6,1}\text{Mon}_{d}+\beta_{h,6,2}\text{Sat}_{d}+\beta_{h,6,3}\text{Sun}_{d}}_{\text{$=\bm{\beta}_{h,6}^{\prime}\mathbf{Cal}_{d}$}}+\nu_{d,h},$ (2)
@@
3.2 nonlinear Networks and Combined structures
The second main architecture in this project is a more elaborate network that, besides the skip connection from the input to the output layer, employs a hidden layer equipped with the Leaky ReLU activation function to capture potential nonlinearities. This activation function is well-suited to electricity price data as it solves the dead-neuron problem that can arise in the presence of negative prices.
Two specifications of this architecture are considered: the MLP with RLin and the MLP with FLin. The MLP with RLin is the sum of the linear predictor of equation (1) and a nonlinear part $\mathrm{NN}(\theta)$:
@@
where $\mathbf{x}\in\mathbb{R}^{m}$ is the input vector with $m$ regressors, $\mathbf{W}^{(1)},\mathbf{W}^{(2)},\mathbf{b}^{(1)},\mathbf{b}^{(2)}$ are the trainable weight matrices and bias vectors, respectively. $\mathbf{\theta}$ denotes the collection of all trainable parameters and $\alpha=0.01$ represents the default leakage factor in the Leaky ReLU activation function.
Figure 3 illustrates the architecture of MLP with RLin Model. The green neurons represent the input layer. The hidden layer is defined by the blue $n$ neurons, where the number of neurons $n$ is a hyperparameter to be tuned. The output layer consists of 24 neurons, where each neuron provides the forecast for a specific hour of the day, is highlighted in orange. The blue edges linking the input and hidden layers correspond to the weight matrix $\mathbf{W}^{(1)}$, while the remaining blue connections are saved in $\mathbf{W}^{(2)}$. For simplicity, the bias vectors $\mathbf{b}^{(1)}$ and $\mathbf{b}^{(2)}$ are not shown in Figure 3. Note that the rectangular neurons represent vectors rather than scalars. Thus, any connection from a rectangular node to another neuron does not represent a single weight, but a vector of weights. For instance, the blue edges connecting $\mathbf{Com}_{d-2}$ to the hidden neuron $\mathbf{H}_{1}$ illustrate four learnable weights, since $\mathbf{Com}_{d-2}$ is a 4-dimensional vector.
Figure 3: MLP with RLin model architecture for forecasting hourly electricity prices. The red connections represent the skip (linear) connections, while the blue connections correspond to the MLP component. Rectangular neurons denote the vector input.
The red edges directly linking the input to the output, referred to as skip connection, represents the model’s linear predictor. The red solid lines connect to the variables that vary for each hour, whereas the dashed red lines correspond to variables that are constant throughout the day, such as fundamentals. However, an exception arises for the covariate $P_{d-1,23}$, which is constant across hours and also serves as the lagged price for predicting $P_{d,23}$. Accordingly, the skip connection for hours $P_{d,0}$ through $P_{d,22}$ includes 15 regressors, while the skip connection for $P_{d,23}$ contains only 14 covariates. It is worth noting that the linear relationships are not only learned through skip connection since the MLP component of the model can learn linear dependencies as well, which results in identifiability issues. To address this, one could incorporate the orthogonalization cell proposed by [rugamer] to restore the interpretability of the linear components. In this study, however, we do not apply this procedure, as our primary focus is on predictive performance rather than on coefficient interpretability.
MLP with FLin model shares the same hidden layer design as the MLP with RLin model, which is defined as $\mathrm{NN}(\mathbf{\theta})$ in equation (3). Yet differs in how the input-to-output connection is structured, as this linear component corresponds to that of the FLin model. The model equation is expressed in the following form:
@@
Hence, the MLP with FLin model employs 175 regressors to forecast each hour.
In addition to the four models introduced above, we develop three further variants. The first is the plain MLP model, which consists of an input layer, a hidden layer, and an output layer, without any skip connections such that the coefficients of the linear component in equation (3) are set to zero ($\beta_{h,0}=\beta_{h,1}=\beta_{h,3}=\mathbf{\beta}_{h,4}^{\prime}=\mathbf{\beta}_{h,5}^{\prime}=\mathbf{\beta}_{h,6}^{\prime}=0$). The second and third are the RLin with OLS and the MLP with RLin and OLS models. These are structurally similar to the RLin and the MLP with RLin models, respectively, but differ in their weight initialization: the OLS-based variants use the coefficients of a simple linear model estimated via OLS as initial weights, whereas the non-OLS counterparts initialize weights randomly.
It is important to highlight that each of the above models is implemented identically to predict electricity prices in the German-Luxembourg and Spain bidding zones. The only exception is that $\beta_{h,4,3}$ is set to zero in the Spanish case due to the absence of offshore wind generation data. Furthermore, all the variables are standardized to enhance training stability. For each training window, standardization is performed by subtracting the mean and dividing by the standard deviation, employing the window-specific mean and standard deviation. The same statistics are subsequently used to standardize the observation to be forecasted with respect to the corresponding training window. This design ensures that the transformation relies only on information available within that window and that no future information is utilized, thereby preventing data leakage.